During routine​ conversations, the CEO of a new​ start-up reports that 15​% of adults between the ages of 21 and 39 will purchase her new product. Hearing​ this, some investors decide to conduct a​ large-scale study, hoping to estimate the proportion to within 5​% with 90​% confidence. How many randomly selected adults between the ages of 21 and 39 must they​ survey?

The number of adults that should be surveyed is ------. ​(Round up to the nearest whole​ number.)
Answer #1

Given : Margin of error=E=5%=0.05

Significance level=0.10

Estimate of the sample proportion=\hat{p}=15%=0.15

Now , the critical value is , Z_{\alpha/2}=Z_{0.10/2}=1.645 ; From Z-table

Therefore , the required sample size is ,

n=\frac{Z_{\alpha/2}^2*\hat{p}(1-\hat{p})}{E^2}=\frac{1.645^2*0.15(1-0.15)}{0.05^2}\approx 138

Therefore , The number of adults that should be surveyed is 138.